# Polylog-time- and constant-space-overhead fault-tolerant quantum computation with quantum low-density parity-check codes

### v1

@@ VI.3 Proof of existence of threshold
$p^{\prime}_{j}\leq\omega 2^{2^{d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}},$ (352)
@@
$p_{\mathrm{phys}}$ $\leq\omega 2^{2^{d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}.$ (354)
@@
$\omega 2^{2^{d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{2d_{0}}}<p_{\mathrm{dec}}^{\mathrm{th}},$ (356)
@@ VII Conclusion
Note on a related work
After our work in Refs. [1, 2] had been out, personal communication [86], let us know that their upcoming work using substantially different techniques from ours will also provide an analysis of fault-tolerant quantum computation with polylogarithmic time overhead and constant space overhead.
@@ A.4 Threshold theorem for open circuits
Starting from $W^{(L)}\leq c_{1}W(C)$ where $c_{1}$ is constant, we obtain
@@
where $d_{2}$ accounts for additional depth due to the decoding gadget. For simplicity, we can bound $D^{(l-1)}\leq w_{2}D^{(l)}$ with $w_{2}=d_{1}+d_{2}$. Starting from $D^{(L)}\leq c_{2}D(C)$ where $c_{2}$ is a constant, we obtain
@@
The failure probability of the reduction is bounded, due to the union bound, by
$q_{j}\times|C^{(L)}|$ $\leq q_{j}\times 4W(C)\times 4D(C),$ (428)
@@
where we used (416) and $c_{3}\geq 16$ is a positive constant. Thus, a pair $(\tilde{C}(C,\delta_{3}),\{p_{i}\}_{i\in\tilde{C}(C,\delta_{3})})$ where $p_{i}\leq p_{\mathrm{loc}}$ for all $i\in\tilde{C}(C,\delta_{3})$ is $\delta_{3}$-reducible to another pair $\quantity(C\cup C^{\mathrm{wait}},\{q_{j}\}_{j\in C\cup C^{\mathrm{wait}}})$ satisfying (426) and (427), where we use the equivalence of $C$ and $C^{(L)}$ in the sense in Def. 12.

### v2

@@ VI.3 Proof of existence of threshold
$p^{\prime}_{j}\leq 2^{2^{d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{d_{0}}}$ (365)
@@
$p_{\mathrm{phys}}$ $\leq 2^{2^{d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{d_{0}}}.$ (367)
@@
$2^{2^{d_{0}}}\quantity(1+\frac{dp_{\mathrm{loc}}}{\lambda})^{2^{d_{0}+1}(d_{0}+1)}{\lambda}^{1/2^{d_{0}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+2^{2^{d_{0}+1}(d_{0}+1)}(dp_{\mathrm{loc}})^{1/2^{d_{0}}}<p_{\mathrm{dec}}^{\mathrm{th}},$ (369)
@@ VII Conclusion
Note added
: After the completion of our manuscript, a paper presenting a fault-tolerant protocol with constant space overhead and polylogarithmic time overhead [85] was uploaded, relying on a substantially different approach from ours. The primary contribution of our work, distinct from theirs, lies in the proof of the threshold theorem in Theorem 11, which takes into account the non-zero runtime of classical computation such as that of decoding algorithms as presented in Assumption 3. Furthermore, their work depends on unpublished assertions required for the existence of positive-rate quantum locally testable codes (QLTCs) with inverse-polylogarithmic relative distance, and inverse-polylogarithmic soundness (see private-communication citations in Theorems 6.9 and 6.11 (footnote 40) of Ref. [85]). In contrast, our work completes the proof of the threshold theorem with constant space overhead and polylogarithmic time overhead without relying on conjectures.
@@ A.4 Threshold theorem for open circuits
Since there exists a positive constant $c_{1}$ such that $W^{(L)}\leq c_{1}W(C)$, we obtain
@@
where $d_{2}$ accounts for additional depth due to the decoding gadget. For simplicity, we can bound $D^{(l-1)}\leq w_{2}D^{(l)}$ with $w_{2}=d_{1}+d_{2}$. Since there exists a positive constant $c_{2}$ such that $D^{(L)}\leq c_{2}D(C)$, we obtain
@@
The failure probability of the reduction is bounded by the union bound. Since there exists a positive constant $c_{3}$ such that $|C^{(L)}|\leq c_{3}W(C)D(C)$, we have
$q_{j}\times|C^{(L)}|$ $\leq q_{j}\times c_{3}W(C)D(C),$ (441)
@@
where we used (429). Thus, a pair $(\tilde{C}(C,\delta_{3}),\{p_{i}\}_{i\in\tilde{C}(C,\delta_{3})})$ where $p_{i}\leq p_{\mathrm{loc}}$ for all $i\in\tilde{C}(C,\delta_{3})$ is $\delta_{3}$-reducible to another pair $\quantity(C\cup C^{\mathrm{wait}},\{q_{j}\}_{j\in C\cup C^{\mathrm{wait}}})$ satisfying (439) and (440), where we use the equivalence of $C$ and $C^{(L)}$ in the sense in Def. 12.